# TurboMap: GPU-Accelerated Local Mapping for Visual SLAM

### v1

@@ Abstract
This paper presents TurboMap, a GPU-accelerated and CPU-optimized local mapping module for visual SLAM systems. We identify key performance bottlenecks in the local mapping process for visual SLAM and address them through targeted GPU and CPU optimizations. Specifically, we offload map point triangulation and fusion to the GPU, accelerate redundant keyframe culling on the CPU, and integrate a GPU-accelerated solver to speed up local bundle adjustment. Our implementation is built on top of ORB-SLAM3 and leverages CUDA for GPU programming. The experimental results show that TurboMap achieves an average speedup of $1.3\times$ in the EuRoC dataset and $1.6\times$ in the TUM-VI dataset in the local mapping module, on both desktop and embedded platforms, while maintaining the accuracy of the original system.
@@ I Introduction
A typical visual SLAM pipeline consists of three main components: tracking, local mapping, and loop closing [1]. Tracking estimates the camera pose by extracting features from the current frame and matching them to existing map points, local mapping refines and expands the map to improve accuracy of the system, and loop closing detects revisited locations to correct accumulated drift. Well-known examples of this class of SLAM systems include ORB-SLAM [2, 3, 4], OpenVINS [5], and Kimera [6].
In real-time visual SLAM, each component must operate within strict time constraints, as delays in any stage compromise overall system performance, with local mapping being particularly critical. In systems that maintain a global map, such as ORB-SLAM, falling behind or skipping local mapping results in larger, less accurate maps and a higher risk of tracking loss. Systems without a global map (e.g., Kimera, OpenVINS) avoid immediate tracking failures but experience delayed or missed loop closures. However, because these systems still rely on optimized maps for accurate pose estimation, incomplete local mapping ultimately degrades trajectory quality [1].
To avoid keyframe drops or processing delays in local mapping, which can degrade localization accuracy, incoming keyframes generated by the tracking module must be processed promptly. To address this, we present TurboMap, a GPU-accelerated and CPU-optimized local mapping module for visual SLAM systems. Our approach begins by analyzing local mapping’s computational bottlenecks. We then accelerate these bottlenecks either by offloading their logic to the GPU or by optimizing them on the CPU. For GPU-based tasks, we carefully design kernels and memory management strategies to minimize overhead, such as kernel launch costs and data transfers, ensuring that GPU parallelization yields performance gains.
@@
We compare the local mapping performance of TurboMap with the vanilla ORB-SLAM3 using widely used datasets EuRoC [7] and TUM-VI [8]. Our experiments show that TurboMap achieves an average speedup of $1.3\times$ on the EuRoC dataset and $1.6\times$ on the TUM-VI dataset in the local mapping module compared to ORB-SLAM3, across both desktop and embedded platforms, while maintaining similar trajectory accuracy. Although our implementation is based on ORB-SLAM3, we expect the proposed techniques to generalize to other visual SLAM systems that maintain a global map and implement the components we accelerate in TurboMap. Our code is available at https://github.com/sfu-rsl/TurboMap.
@@ II Related Work
In this section, we briefly discuss related work on GPU acceleration for tracking, local mapping, and loop closing modules in the context of ORB-SLAM based systems [2, 3, 4].
@@
Other works focus on accelerating least squares subproblems in various modules. Gopinath et al. [12] parallelize matrix calculations to accelerate local bundle adjustment in local mapping, while Kumar et al. [13] offload Jacobian computations for pose graph optimization to the GPU. The Ceres Solver [14] introduces optional CUDA-based linear solvers to speed up the optimization steps in bundle adjustment, and MegBA [15] provides a fully GPU-native, distributed BA framework capable of handling large-scale optimizations.
In contrast to previous work, we focus on offloading local mapping tasks aside from bundle adjustment to the GPU. Specifically, we move the logic of map point creation and fusion to the GPU and accelerate redundant keyframe culling on CPU. To support these parallelized GPU algorithms, we also design a dedicated keyframe storage system that efficiently manages the data required by these operations.
@@ III Background
The local mapping module in a visual SLAM system is responsible for processing keyframes as well as building, maintaining, and optimizing the map. The computations performed in this component are critical for ensuring both the efficiency and accuracy of the system. Figure 1 illustrates the different stages of local mapping, using ORB-SLAM3 as an example visual SLAM system. In the rest of the section, we explain each component.
@@ III-A Map-point Creation
Map point creation is responsible for estimating the 3D coordinates of 2D keypoints and integrating the resulting 3D points into the map. The system begins by triangulating each keypoint in the current keyframe with the corresponding keypoints in neighboring keyframes to determine their 3D positions. After triangulation, a set of geometric and photometric constraints is evaluated, and new map points are inserted into the map only if these constraints are satisfied. The constraints include sufficient parallax between views, positive depth (i.e. points must lie in front of the camera), low re-projection error, and consistency in scale.
@@ IV Design
To reduce the latency of the local mapping process, we focus on optimizing its most time-consuming components. As shown in Figure 2, approximately 95% of the computation time is spent on Map Point Creation, Map Point Fusion, LBA, and Keyframe Culling.
Figure 3 illustrates the local mapping architecture in TurboMap, showing the division of responsibilities between the CPU and the GPU. We offload a component to the GPU if it meets three key criteria: It has a high computational overhead, exhibits a high degree of parallelism, and requires only a small amount of input data. If any of these conditions are not satisfied, GPU acceleration may offer limited benefits or even introduce additional overhead due to data transfer and synchronization costs. To further enhance performance, we allocate GPU memory during system initialization and release it upon shutdown, minimizing runtime allocation costs. We discuss the different components used in TurboMap in the following sections.
IV-A System Overview
To explain how TurboMap operates and to highlight its differences and advantages over the original ORB-SLAM3, we present the local mapping flow depicted in Figure 3 using ORB-SLAM3 as the baseline for comparison. The local mapping stage in both systems is triggered when tracking hands a newly created keyframe to it. In ORB-SLAM3 all keyframes are maintained in CPU memory. In contrast, as shown in Figure 3, TurboMap immediately copies each new keyframe to a dedicated GPU storage, enabling subsequent components to access required data without repeated transfers.
The first operation in the pipeline is to remove recently created map points that are redundant or classified as outliers. This check is inexpensive and remains on the CPU with its logic unchanged. The next stage is Map Point Creation, which involves searching for corresponding map points in neighboring keyframes and then performing triangulation. In ORB-SLAM3, both steps are executed sequentially on the CPU, with the search phase dominating the computational cost of this component. In TurboMap we separate these two sub-tasks and offload the correspondence search to the GPU while retaining triangulation on the CPU. Triangulation has negligible per-match cost and would incur unnecessary data movement if performed on the GPU, so keeping triangulation on the CPU provides the best trade-off between computation and transfer overhead. On the other hand, the correspondence search stage is computationally heavy and highly parallelizable, so TurboMap offloads it to the GPU.
After creating new map points, the system performs duplicate map point detection and fusion. Fusion compares candidate descriptors across first-order and selected second-order neighbors to find duplicates, which involves many independent descriptor comparisons. Because fusion is computationally intensive and highly parallelizable, TurboMap offloads it to the GPU, where parallel execution significantly reduces latency compared to the CPU implementation in ORB-SLAM3. The pipeline then performs Local Bundle Adjustment to refine the poses of local keyframes and the positions of associated map points. As one of the primary bottlenecks in local mapping, reducing its latency is essential to achieve faster overall performance. Instead of re-implementing this complex optimizer, TurboMap integrates an existing GPU-accelerated LBA solver [12] and uses it as a drop-in module for ORB-SLAM3’s optimizer. Finally, redundant keyframes are removed by Keyframe Culling. In TurboMap, this stage is redesigned on the CPU rather than ported to the GPU, since it requires frequent access to map point data that are continuously updated during local mapping. Transferring these data to the GPU at each iteration would introduce significant overhead, so TurboMap instead optimizes the CPU implementation to reduce latency.
In the following subsections, we provide a detailed description of each component in TurboMap and the rationale behind its GPU or CPU optimization.
@@ IV-B Keyframe Storage
Search for Triangulation and Map Point Fusion, the two components which TurboMap offloads to the GPU, frequently need access to map point data from many neighboring keyframes. A naive approach would transfer all required keyframes to the GPU each time a new keyframe is processed. However, this is highly inefficient: Each keyframe is roughly 1 MB, and with tens of neighbors, repeated transfers would introduce substantial overhead. To solve this issue, TurboMap implements a persistent keyframe storage on the GPU. Each keyframe is transferred to the keyframe storage on GPU once when it enters local mapping and remains there until it is removed from the map or the system shuts down, avoiding repeated data transfers and ensuring efficient GPU execution. The keyframe data stored in this storage mainly include keypoints and their descriptors for each keyframe.
IV-C Search for Triangulation
As shown in Figure 3, the first component we offload to the GPU is Search for Triangulation. This stage is computationally intensive and highly parallelizable, making it well suited for GPU acceleration with minimal data transfer overhead. During the Search for Triangulation stage, the system attempts to identify the most similar feature in neighboring keyframes for each feature in the current keyframe. This is done by computing the descriptor distance between each pair of features. Additionally, each potential match is validated against epipolar constraints to ensure geometric consistency. Given the high number of features per keyframe and the large number of descriptor comparisons required, this process is computationally expensive. However, since finding the best match for each feature is independent of other features and keyframes, the task is highly parallelizable. To take advantage of this situation, we offload this task to the GPU and task each GPU thread to independently perform descriptor comparisons and verify epipolar constraints. This parallel execution significantly reduces latency, as the workload is distributed across many threads. In terms of the required data, the keyframes represent the largest potential overhead. However, since we use keyframe data that is already stored on the GPU, no additional data transfer is incurred.
@@ IV-D Map-point Fusion
Duplicate Map Point Fusion is the second component which TurboMap offloads to the GPU as shown in Figure 3. The primary goal of this function is to determine whether there exists a map point in the map that (1) has descriptors very similar to those of a map point associated with the current keyframe and (2) is spatially close to that map point. This process is computationally intensive, highly parallelizable, and benefits from the fact that most of the required data, particularly keyframe-related information, already reside in GPU memory.
The Fuse and Search for Triangulation components share similar logic, as both compute the feature similarities. However, as described in Section III, the Fuse component searches for duplicate map points across first-order and selected second-order neighboring keyframes, whereas Search for Triangulation is limited to first-order neighbors. As a result, the fuse operation is more computationally intensive, which is also evident in Figure 2. This makes Fuse even a better candidate for parallelization. In the GPU implementation, each thread is responsible for identifying the most similar map point to a given point in the current keyframe within a specific neighboring keyframe. Regarding data transfer, the primary data required by this component includes the neighboring keyframes and the map points associated with the current keyframe. The former is already stored on the GPU, incurring no transfer cost, while the latter amounts to approximately 100 KB, which is an insignificant size that has minimal impact on the performance of the GPU program.
@@ IV-E Local Bundle Adjustment
The final component that TurboMap accelerates by using GPU computation is Local Bundle Adjustment, as illustrated in Figure 3. Local Bundle Adjustment is the most time-consuming stage in local mapping, accounting for approximately 60% to 75% of the total execution time, as shown in Figure 2. As discussed in Section II, several prior efforts have been made to accelerate this component due to its significant performance impact. Since this component has been extensively studied and optimized in prior work, we chose to adopt an existing implementation rather than develop a new one from scratch. Specifically, we integrate the GPU-accelerated local bundle adjustment from the work of Gopinath et al. [12], which is compatible with ORB-SLAM3. In this work, the Schur Complement method, which is used to reduce the system size and speed up the LBA process, is offloaded to the GPU. Their implementation serves as a drop-in replacement block solver for g2o [16], the optimization library used by ORB-SLAM3 for solving the local bundle adjustment problem.

### v2

@@ Abstract
In real-time Visual SLAM systems, local mapping must operate under strict latency constraints, as delays degrade map quality and increase the risk of tracking failure. GPU parallelization offers a promising way to reduce latency. However, parallelizing local mapping is challenging due to synchronized shared-state updates and the overhead of transferring large map data structures to the GPU. This paper presents TurboMap, a GPU-parallelized and CPU-optimized local mapping backend that holistically addresses these challenges. We restructure Map Point Creation to enable parallel Keypoint Correspondence Search on the GPU, redesign and parallelize Map Point Fusion, optimize Redundant Keyframe Culling on the CPU, and integrate a fast GPU-based Local Bundle Adjustment solver. To minimize data transfer and synchronization costs, we introduce persistent GPU-resident keyframe storage. Experiments on the EuRoC and TUM-VI datasets show average local mapping speedups of $1.3\times$ and $1.6\times$, respectively, while preserving accuracy.
@@ I Introduction
A typical visual SLAM pipeline consists of three main components: tracking, local mapping, and loop closing [1]. Tracking estimates the camera pose by extracting features from the current frame and matching them to existing map points, local mapping refines and expands the map to improve accuracy of the system, and loop closing detects revisited locations to correct accumulated drift. Well-known examples of this class of SLAM systems include ORB-SLAM [2, 3, 4], VINS-Mono [5], and Kimera [6].
In real-time visual SLAM, each component must operate within strict time constraints, since delays in any stage degrade overall system performance, with local mapping being particularly critical. This requirement is even more pronounced in systems that maintain a global map [4] [6] [7]. Such systems generally achieve higher accuracy due to consistent map optimization and loop closure, but suffer from degraded performance in highly concurrent environments due to inefficient concurrent access to the shared global map [8]. As a result, local mapping performance becomes especially important, as falling behind or skipping local mapping leads to less accurate maps and increases the risk of tracking loss.
To mitigate these issues, achieving low latency in local mapping is essential. One promising approach is to parallelize the most computationally intensive tasks, for which GPUs are well suited. However, enabling parallel computation in local mapping is challenging due to the algorithmic structure of visual SLAM systems. First, local mapping frequently reads and modifies shared map structures while other modules are running concurrently and accessing the same data. To ensure consistency, these operations use synchronization (e.g., using locks), forcing many computations to wait for exclusive access and thereby serializing the execution. Consequently, without restructuring the underlying computations, the design of visual SLAM systems inherently limits the effectiveness of parallel execution in local mapping. Second, local mapping operates on large keyframe and map point data structures. Transferring these data to the GPU at each iteration can introduce substantial data movement overhead, which may offset the computational gains of parallel execution if not carefully managed. These challenges explain why prior acceleration efforts [9][10][11] have mainly focused on components that are naturally data-parallel and do not involve shared-state updates or synchronization. To the best of our knowledge, no previous work addresses the local mapping module holistically while explicitly tackling both synchronization-induced serialization and data transfer overhead.
We propose TurboMap, a redesigned and accelerated local mapping module that enables efficient parallel execution by addressing two fundamental challenges: synchronization-induced serialization caused by shared-state updates and the high overhead of transferring large data structures to the GPU. Our contributions are threefold: (1) we analyze data dependencies and runtime latency within the local mapping pipeline to identify the components that dominate execution time and evaluate whether their computations can be parallelized; (2) we restructure selected computations to remove synchronization and sequential dependencies, enabling parallel execution on the GPU, while refining CPU-side procedures when parallelization does not provide measurable benefits; and (3) for the components executed on the GPU, we introduce a persistent keyframe storage system that keeps required keyframe data resident on the GPU throughout the local mapping process and stores only immutable data, reducing data transfer and synchronization overhead to improve overall execution efficiency.
@@
We compare the local mapping performance of TurboMap with the vanilla ORB-SLAM3 using widely used datasets EuRoC [12] and TUM-VI [13]. Our experiments show that TurboMap achieves an average speedup of $1.3\times$ on EuRoC and $1.6\times$ on TUM-VI in the local mapping module compared to ORB-SLAM3, across both desktop and embedded platforms, while maintaining similar trajectory accuracy. Although our implementation is based on ORB-SLAM3, we expect the proposed techniques to generalize to other visual SLAM systems that maintain a global map and implement the components we accelerate in TurboMap. We will open-source TurboMap upon acceptance.
@@ II Related Work
In this section, we discuss related work on GPU acceleration for tracking, local mapping, and loop closing modules in the context of ORB-SLAM based systems [2, 3, 4].
@@
Other works focus on accelerating least squares subproblems in various modules. Gopinath et al. [9] parallelize matrix calculations to accelerate local bundle adjustment in local mapping, while Kumar et al. [17] offload Jacobian computations for pose graph optimization to the GPU. Ceres Solver [11] introduces optional CUDA-based linear solvers to speed up the optimization steps in bundle adjustment, and MegBA [10] provides a fully GPU-native, distributed BA framework capable of handling large-scale optimizations.
In contrast to previous work on local mapping that primarily focuses on bundle adjustment [9][10][11], we target the remaining components that dominate runtime and are constrained by synchronization and shared-state updates. Specifically, we restructure Map Point Creation and parallelize the Keypoint Correspondence Search stage, and also redesign the Map Point Fusion stage to enable parallel execution. We also optimize Redundant Keyframe Culling on the CPU. To support these parallelized GPU stages, we also design a dedicated keyframe storage system that efficiently manages the data required by these operations. We will introduce these components in Section IV.
@@ III Background
The local mapping module in a visual SLAM system is responsible for processing keyframes as well as building, maintaining, and optimizing the map. The computations performed in this component are critical for ensuring both the efficiency and accuracy of the system. Figure 1 illustrates the different stages of local mapping, using ORB-SLAM3 as an example visual SLAM system. In the rest of the section, we explain the components related to our work.
@@ III-A Map-point Creation
Map point creation is responsible for estimating the 3D coordinates of 2D keypoints and integrating the resulting 3D points into the map. The system begins by performing a keypoint correspondence search across neighboring keyframes to identify matching keypoints. Once the correspondences are found, triangulation is used to determine their 3D positions. After triangulation, a set of geometric and photometric constraints is evaluated, and new map points are inserted into the map only if these constraints are satisfied. The constraints include sufficient parallax between views, positive depth (i.e. points must lie in front of the camera), low re-projection error, and consistency in scale.
@@ IV Design
As shown in Figure 2, approximately 95% of computation time in local mapping is spent on Map Point Creation, Map Point Fusion, Local Bundle Adjustment (LBA), and Keyframe Culling. For GPU-based parallelization, Map Point Creation, Map Point Fusion, and Keyframe Culling are challenging because they frequently update shared data structures, such as map points and keyframes, which requires synchronization to maintain consistency. These shared-state updates create serialization points and limit the benefits of GPU execution. In contrast, previous work has successfully accelerated LBA on the GPU because it is naturally parallelizable and involves minimal synchronization or shared-state updates. In addition to parallelization challenges, all of these components operate on large collections of keyframe and map point data, and repeatedly transferring such heavy data structures to the GPU can introduce significant data movement overhead, further limiting performance if not managed carefully.
To address these challenges, TurboMap takes a holistic approach: we restructure computations, manage data efficiently, and apply parallel execution where feasible. Components are parallelized only if they incur high computational cost, allow the elimination of synchronization that would otherwise hinder parallel execution, and avoid repeated heavy data transfers between the CPU and GPU. In the following sections, we describe each component of TurboMap in detail, explaining the underlying challenges and the corresponding solutions that enable efficient execution. Throughout this discussion, we refer to Figure 3, which provides an overview of the local mapping architecture and illustrates the division of responsibilities between the CPU and GPU.
IV-A Parallelization Challenges
Parallelizing local mapping is constrained by two fundamental architectural challenges. First, the main stages in local mapping are tightly coupled with updates to shared map structures. During Map Point Creation, newly triangulated points are inserted into the global map and linked to observing keyframes. In Map Point Fusion, duplicate points are merged and their observations are reassigned. In Keyframe Culling, removing a redundant keyframe requires updating covisibility relationships and observation lists. In each of these stages, the values of keyframes or map points are modified, and these data structures are shared with other concurrent processes such as tracking and loop closing. Because multiple processes may access or modify the same objects, synchronization is necessary to preserve consistency. As a result, these interleaved updates create serialization points that limit parallel execution.
Second, local mapping operates on large and complex CPU-resident data structures, including keyframes and map points. Each keyframe contains hundreds or thousands of feature descriptors, geometric information, and observation links, while the map may consist of tens of thousands of points. Transferring such large amounts of data to the GPU at each iteration would require substantial bandwidth and memory operations, introducing significant runtime overhead. Moreover, since these structures are frequently updated, naive copying of data would not only be slow but also require careful synchronization to maintain correctness. This makes the size and mutability of the data a major barrier to efficient GPU-based parallelization.
We address the first challenge by restructuring the computations in Map Point Creation and moving the Keypoint Correspondence Search subroutine to the GPU, as well as restructuring the computations in the Fuse stage. To address the second challenge, we introduce a persistent keyframe storage on the GPU. In the following sections, we first describe the keyframe storage design, and then detail how each component addresses the challenge of synchronization-induced serialization.
@@ IV-B Keyframe Storage
Keypoint Correspondence Search and Map Point Fusion, the two components we parallelize on the GPU, frequently require access to map point data from multiple neighboring keyframes. A naive approach would transfer all required keyframes to the GPU whenever a new keyframe is processed. Given that each keyframe occupies approximately 1 MB and that tens of neighboring keyframes may be involved, repeated transfers would introduce significant overhead.
To address this, TurboMap introduces a persistent GPU keyframe storage. We copy each keyframe to the GPU only once when it enters local mapping and keep it resident until it is removed from the map or the system shuts down. To reduce both data transfer and synchronization overhead, we partition keyframe data into immutable and mutable components, storing only the immutable data on the GPU. Immutable data, such as keypoints and their descriptors, do not change after keyframe creation, so keeping them on the GPU eliminates the need for costly synchronization with the CPU. Mutable data, by contrast, can change during local mapping and would require frequent updates if kept on the GPU, which would severely reduce efficiency. We also exclude other keyframe data that are unnecessary for GPU computations. By keeping only the necessary immutable data on the GPU, we avoid repeated transfers and synchronization, and enable efficient parallel execution of parallelized stages.
IV-C Keypoint Correspondence Search
As shown in Figure 3, the first component we parallelize is the Keypoint Correspondence Search stage within Map Point Creation. This stage performs keypoint correspondence search between the current keyframe and its neighboring keyframes. For each keypoint in the current keyframe, the system identifies candidate matches in neighboring keyframes by computing descriptor distances and validating them using epipolar constraints. In the original non-parallel structure, neighboring keyframes are processed sequentially, and triangulation is performed immediately after each valid match is found. Because triangulation inserts new map points into the shared map and updates keyframe and map point structures, these operations require synchronization to preserve consistency. As a result, correspondence search and map updates are tightly interleaved, creating serialization points that prevent parallel execution and make GPU parallelization ineffective.
To address this limitation, we restructure the computation to enable parallel execution. We decouple Keypoint Correspondence Search from the Triangulation step, allowing matches across all neighboring keyframes to be identified before creating new map points and removing synchronization-dependent operations from the search stage. To handle data transfers efficiently, we take two measures. First, we eliminate repeated synchronization by batching the required map point data and transferring it to the GPU in a single step prior to parallel execution. Second, most of the data required at this stage originate from neighboring keyframes. Because these keyframes are already stored in persistent GPU memory as part of our keyframe storage design, no additional transfers are necessary. Each GPU thread then independently performs descriptor comparisons and epipolar validation, distributing the workload across thousands of threads. By combining structural code changes with GPU-resident data, we significantly reduce latency and enable efficient parallel execution.
@@ IV-D Map-point Fusion
Duplicate Map Point Fusion is the second component TurboMap parallelizes on the GPU as shown in Figure 3. The primary goal of this stage is to determine whether there exists a map point in the map that (1) has descriptors very similar to those of a map point associated with the current keyframe and (2) is spatially close to that map point. This process is computationally intensive and has the potential to be parallelizable. It also benefits from our keyframe storage design, as most of the required data, particularly keyframe-related information, already reside in GPU memory. However, similar to Keypoint Correspondence Search, the original structure is sequential, where the system loops over neighboring keyframes and their associated map points, and once a duplicate is found, it immediately fuses the map points. Since fusion modifies map point structures and requires synchronization, this tightly coupled logic prevents straightforward parallelization.
To enable GPU acceleration, we apply a restructuring strategy similar to that used in Keypoint Correspondence Search. Instead of fusing duplicates as they are found, we first determine all candidate duplicate pairs in parallel and defer the actual fusion to a subsequent sequential stage. In the GPU implementation, each thread is responsible for identifying the most similar map point to a given point in the current keyframe within a specific neighboring keyframe. To eliminate synchronization during the search phase, we gather all required map point data in a single batch and transfer it to the GPU for processing. The neighboring keyframe data are already stored on the GPU, incurring no transfer cost, while the map points associated with the current keyframe amount to approximately 100 KB, which has negligible impact on performance. By separating search from fusion and removing synchronization from the parallel stage, we enable efficient GPU execution while preserving the correctness of the original logic.
@@ IV-E Local Bundle Adjustment
The final component that TurboMap accelerates by using GPU computation is Local Bundle Adjustment, as illustrated in Figure 3. As discussed in Section II, several prior efforts have been made to accelerate this component. Since this component has been extensively studied and optimized in prior work, we chose to adopt an existing implementation rather than develop a new one from scratch. Specifically, we integrate the GPU-accelerated local bundle adjustment from the work of Gopinath et al. [9], which is compatible with ORB-SLAM3. In this work, the Schur Complement method, which is used to reduce the system size and speed up the LBA process, is parallelized on the GPU. Their implementation serves as a drop-in replacement block solver for g2o [18], the optimization library used by ORB-SLAM3 to solve the local bundle adjustment problem. The GPU solver has been numerically validated against the original CPU-based g2o implementation, showing negligible differences in reprojection error, estimated states, and convergence behavior. It preserves the same mathematical formulation and optimization framework, ensuring that the acceleration improves efficiency without affecting accuracy or stability.